Multiply:
3s^2 * 5s^7
Is the answer:
15s^14?? Thank you so much!

Answer by checkley75(3666)

3S^2*5S^7=15S^9
WHEN YOU MULTIPLY EXPONENTS YOU NEED TO ADD THEM NOT MULTIPLY THEM.
YOU MULTIPLY THEM WHEN THE EXPONENT IS RAISED BY ANOTHER EXPONENT THUS:
(4^3)^3=4^9 OR (4^3)(4^3)(4^3)=4^9